Now I have and ult=o= o. d 2 +2 I How can I get the eq (x-xo + I) ² - t² = 7 !
Integral equation
Answer #1

Giveni du = (1-4²) 12 ao and u (t=0) = 0 dt du - adt I using variable separable -u232 on equation of Integrating both sides d: Sidt=t and c= Julegration 7 constant put in abo we equation u = got tc Ft Now at tao thao 0 = ao(o)tc To=c] put in u = gotderivate both sides Now let 11790²t2 = ÞI des of ar at dt = dp 290²t dt = dp at = dp 2007 Thence I atet - r got db Ita²42 arfut in to to Xo = a tc (16 2 2 = C = 1 51 +92 +2 + x - 720 ( - Xo +1 = 1/ +4 () savaring both sides 4x=x0+ 1) = 02 (1+90 € (X